answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting a pair of red kings from a regular deck of cards without jokers?

1 in 2652. The odds of drawing one red king are 1 in 52. The odds of drawing the second are 1 in 51. Multiply 52 by 51, and you get 2652.

1/676

How many even numbers in a regular deck of cards?

there's 4 suits in one deck of cards.

if j =11, queen =12, and king =13.

that means there will be 6 even numbers card in one suit, and times that by 4, which mean 24. theres 24 cards are even in a regular deck of cards.

How many red spades in a deck of cards?

In a standard deck of cards, all the spades are black.

What is the probability you roll a prime number on a six sided dice?

it might be 395,652,150,468,457,125,279,456,000,000 chanch to roll a prime number when a monkey appeares at the door

When you toss a coin 5 times probability of getting exactly 4 heads?

Because there are only 2 outcomes for the flip of a coin, for 5 flips you just need to take (1/2)5, which equals 1/32.

This implies there are 32 different outcomes for the case of tossing a coin 5 times. From these 32 outcomes 5 have exactly 4 heads: THHHH, HTHHH, HHTHH, HHHTH, and HHHHT. So the probability of getting exactly 4 heads when you toss a coin 5 times is:

P(4H,!T) = 5/32 = 0.15625 ≈ 15.6%

What is the possible phenotype of four offspring?

It dependes on the genes of the parent if you have no information of the genes of the parent I cannot tell you

Why is it called experimental probability?

It's difficult to think of a real event to which an exact probability can be assigned. We say that flipping a coin yields 'heads' with probability 1/2 but we do not know that definitely. The only way of assigning a probability in the sense of numbers of heads versus total numbers of flips is by experiment. (Be aware though that there are other interpretations of the word probability.) If I were to flip a coin 500 times and obtained 249 heads then the experimental probability of obtaining a head would be 249/500 or 0.498.

Where is probability found?

Wherever there is some random (that is, non-determinate) element in the occurrence of an event.

What is the first step in finding the probability of getting two numbers whose sum 2?

The first step is to determine what the experiment is: rolling a pair of dice, drawing two cards from a deck of playing cards, selecting numbered cards?

How does the probability improve of choosing the one correct answer as incorrect answers are eliminated one at a time from 5 multiple choice question such as those found on the SAT?

Probability of hitting the bull's-eye on the next random shot after eliminating . . .

- no answer . . . . . 20%

- 1 answer . . . . . . 25%

- 2 answers . . . . . 331/3%

- 3 answers . . . . . 50%

- 4 answers . . . . . 100%

- 5 answers . . . . . zero

Probability of choosing the one correct answer increases significantly from

these figures if you actually know something about the subject.

What is a probability of rollong an even number on a number diece?

A die has 6 sides with the numbers 1,2,3,4,5 and 6. Three of those numbers are even. So the probability of rolling an even number is 3 out of 6, or 1 out of 2.

What is the probability the cube will land on 4 and the coin will land heads up?

There is a 1/6 chance of rolling a 4 on a fair die, and a 1/2 chance of a fair coin landing heads up. Multiply 1/6 X 1/2. The probability of both happening is 1/12.

Why do you marry someone?

Because you can't live with out them not being married to you.

What is a frequency distribution defined by its average and standard deviation?

The triangular, uniform, binomial, Poisson, geometric, exponential and Gaussian distributions are some that can be so defined. In fact, the Poisson and exponential need only the mean.

8 dices are thrown probability of even numbers?

Since there are only odd & even numbers, the probability is 1/2 or 0.5.

A single die is rolled what is the probability tha the number rolled will be greater than 3?

Total different outcomes = 6

Successful outcomes = 3 (rolls of 4, 5, or 6)

Probability of success = 3/6 = 1/2 = 50%

What does the probability density function tell you?

The area under the pdf between two values is the probability that the random variable lies between those two values.

What is the probability of independent events?

If you mean the probability that independent events will occur, it is the product of the individual probabilities that they occur. Ex: Probability of picking a red marble from a jar with 4 red and 6 green marbles and picking a black card from a deck of 52.

P(red) = 4/10 = 2/5

P(black card0 = 26/52 = 1/2

P(both) = 2/5 x 1/2 = 1/5

Trending Questions
How many kings in deck of cards? How many colors would be needed to color a map of US states without any state touching the same colors? What does chance of rain mean? If you flip four quarters all at once what is the likelihood that you will get 2 heads and 2 tails? What is marginal probability? If person A can hit a target 6 times in 7 shots and B can hit it 4 times in 5 and C can hit it 3 times in 4 then what is the probability the target will be hit exactly twice if A B and C take one shot? What is the probability for choosing a red marble? What does the expression mutually exclusive mean? 4 dice each dice has 6 numbers how many combinations could there be? What is the probability of drawing 5 face cards from a standard deck of cards? What is an example of how a standard cost is established? What is the safest searching process is to search Before entering an intersection? Identify any means of distribution? What is the probability of rolling a 0 on a two six sided dice? What is the probablity of rolling a number less than or equal to 8 with two dice given that at least one of the dice must show a 6? What is the probability that the first person selected is a supervisor and in management The table of information is here tinyurl.comc4eqpx? How do you roll a big number when playing dice? What is the probability of picking an s from HOUSE? What is the probability of not being dealt a queen in a 52 card deck? What is the difference between probability and statistics?